Subject: APO 3.5m users committee minutes, 3/4/13 meeting
From: Suzanne Hawley
Submitted: Sun, 10 Mar 2013 23:23:11 -0700 (PDT)
Message number: 1306
(previous: 1305,
next: 1307
up: Index)
APO 3.5-m Users Committee Phonecon, 3/4/13
Attending: Mark Klaene, Michael Strauss, John Stocke (for John Bally),
Jon Holtzman, Mary Beth Kaiser, Bill Ketzeback, Joe Huehnerhoff, and Bruce
Gillespie.
Absent: Suzanne Hawley, Scott Anderson, Al Harper, Remy Indebetouw
**********************************
User feedback and comments from institutional representatives:
o Washington (Scott Anderson) - no report.
o Chicago (Al Harper) - no report.
o New Mexico State (Jon Holtzman) - Jon had nothing new to report.
o Princeton (Michael Strauss) - Michael had nothing new to report, other
than everything is going well and the PU users are contented.
o Colorado (John Stock) - John mentioned that there is field trip of CU
students planned in April, and gave thanks to the site staff for their
support in arranging and hosting the visit.
o Virginia (Remy Indebetouw) - no report.
o Johns Hopkins (Mary Beth Kaiser) - Mary Beth said that there is interest
from the JHU users in the possibility of a new optical imager.
**********************************
Discussion of telescope/instruments report:
Mark said discussions have resumed with the US Air Force regarding light
pollution from Holloman AFB, this time initiated by them because of their
interest in energy conservation. Mark also highlighted the DIS binning
issue, where it has been determined that binning odd numbers of pixels in
both (not just one) directions causes binning to fail -- this is
considered a relatively minor problem that would be very difficult to fix,
so we are not planning to work on it any time soon. Work continued on the
echelle guider camera shutter; it is working now and will be replaced
soon. For NIC-FPS, NFS > 8 is still not working and we are no longer
actively trying to fix it as we have run out of options to try.
Mark Klaene's detailed report for the reporting period follows:
*************************
3.5-m Telescope and Instruments Highlights
1/30/13 through 2/27/13
0) Overview
Some weather-associated lost time. Frequent storm systems, about once a
week, continue. Unfortunately, despite the clouds, snowfall accumulation
continues well below average with each storm only depositing a couple of
inches. No clear time was lost due to snow and ice on the roof/lightning
arrestors. Discussions have started with a DOD contractor regarding
reducing light pollution from HAFB.
1) Telescope
The telescope performance has been nominal this period. No further
problems with M2 and M3 actuators have surfaced, but they are still being
monitored closely.
2) Instruments
DIS has been operating nominally. Both red and blue cameras have been
functioning normally with good vacuum and temperature. Further
characterization of the binning issues revealed that binning fails when
both x and y directions are odd number of pixels.
Agile had no problems this period.
SPIcam's new ion pump controller (Gamma vacuum) was upgraded to handle
vacuum spikes better.
TripleSpec operated with no problems. Users should be aware that the slit
is still locked in the 1.1" position.
Echelle servicing was completed and it is operating normally. The problem
of the ECam images showing streaks was traced to a sticking shutter. The
failing shutter was worked on and a new replacement has been ordered. We
are also working with Apogee Instruments on installing their new 5-million
cycle shutter this summer.
NIC-FPS is operating in shared-risk observing mode due to occasional image
corruption issues. NFS greater than 8 still fails. We have exhausted all
ideas to correct this. A glow that has been reported has not been
reproduced, despite numerous attempts on cloudy nights. Further tests are
planned.
**********************************
GIFS training:
Mark reported that the GIFS instrument is working, as are the basic
software actors. The TUI integration will occur after the summer
shutdown. We are canceling the GIFS training session planned for May, and
will try again in the fall.
**********************************
Summer shutdown plans:
Bill said that the summer shutdown duration will be about four weeks. It
is undecided as of yet whether we will break the shutdown into two blocks,
or do it all contiguously. The shutdown work list includes
realuminization of the secondary and tertiary mirrors, servicing their
motion actuators, in situ washing of the primary mirror, annual instrument
servicing (including installation of a new DIS red field lens), inspection
of the azimuth drive motors, and the migration of the TCC software from
VMS to Unix.
**********************************
Instrument documentation:
Jon Holtzman feels that it is important to clean up, update, and generally
improve our user documentation as we transition to new ARC partners.
Although our user documentation is better than it used to be, descriptions
of instrument performance and quirks are not well covered. John Stocke
said that we should identify user documentation "point" persons at our
institutions, and Jon Holtzman agreed that it would be useful if the Users
Committee members find users at their respective institutions to review
and critique the telescope and instrument documentation, which led to a
new action item.
===> ACTION: Users Committee members to find users at their respective
institutions to review and critique the telescope and instrument
documentation.
**********************************
DIS slitviewer camera:
Mark asked the committee members if there was any feedback regarding the
proposal to upgrade the DIS slit viewer, discussed at length at our last
meeting. There was none. Bill said that based on comments he has
previously received on the proposal, we are tweaking the designs and
moving toward PDR. He is looking at the possibility of filters in the
ssystem. Mark asked the committee to inform him of any users who might be
interested in participating in the upcoming review.
**********************************
New APO imager:
Joe and Bill addressed comments and questions on the white paper that was
sent the week before the meeting to the users committee members,
describing requirements and designs for a new optical imaging camera to
replace SPICam. John Stocke noted that the imaging niche of our telescope
is narrow-band, medium-field imaging, in his opinion. The driver to go to
a larger field in a new imager is not compelling; if we can reuse our
existing filters with a focal reducer, that would be ideal. Jon added
that money is a big factor, and the question boils down to whether we can
afford a new instrument that simply maintains our existing capabilities,
albeit with some improvements and maintainability, or do we want to expand
the instrument capabilities, e.g., a wider field, at the same time. John
Stocke repeated that more field of view is not as high a scientific
priority, nor unique to mid-sized telescope. He advocates simultaneous
multi-band imaging, and higher image quality, e.g., tip-tilt correction,
as higher priorities. Mark said that we are severely cost-constrained on
where to take this project -- we encourage further discussion and users
should send us comments or questions. We should talk more on this next
month, especially the scientific case for a 12-arcminute vs. 8-arcminute
field of view for the new imager. Jon Holtzman added that we also need to
discuss photometric and baffling requirements, and whether the instrument
should be mounted at NA2 or on a side port (like Agile). The users
committee was encouraged to disseminate the white paper to interested
users and gather feedback for the next meeting. This is a new action
item.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Send new ARC partner suggestions to Suzanne and Rene.
Status: Open, committee members need to keep sending suggestions.
We need the names of institutions (of any size) and a named contact.
===> ACTION: Users committee members should discuss APOGEE fiber feed to
3.5-m proposal with their users and forward feedback, and ideally name(s)
of interested users who would be willing to participate in an ATI proposal
to Suzanne, Mark and Fred Hearty.
Status: Closed for now, reopen after the APOGEE MRI proposal
request for a 2nd instrument is decided.
===> ACTION: Committee members to talk with their users re LSST followup
programs, and discuss at next Users Committee meeting.
Status: Open, not discussed. We need to talk about this at the
next meeting to provide feedback for Rene Walterbos, who will attend an
NOAO meeting on this subject in mid-April.
===> ACTION: All users should send comments to Suzanne Mark and Bill
about the DIS slitviewer upgrade.
Status: Open, comments still welcome.
[new from this meeting]:
===> ACTION: Users Committee members to identify users at their
respective institutions to review and critique the existing telescope and
instrument documentation.
===> ACTION: Users Committee members to disseminate new APO imager white
paper and gather feedback from their users for discussion at next meeting.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 8 April
at 8:30 AM Pacific Time. The agenda and other materials will be sent to
the committee members during the preceding week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1306 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO